Calculate the following in 6.79 g Li3PO4 a) Li+ ions b) grams O
1
Expert's answer
2016-11-08T14:54:09-0500
a) The mass of Li+ is = 3*7*6.79 g/(3*7 + 31 + 64) = 1.23 g
b) The mass of O is = 4*16*6.79 g/(3*7 + 31 + 64) = 3.75 g